A line has a slope of 3 and contains the point (4, 6). Which is the equation of the line. y=3x-12, y=3x-6, y=3x-14 or y=3x+2?
Answer by kvlsravya(8)

Equation of a line with slope 'm' and passing through the point (x1,y1) is
y - y1 = m(x - x1)
From the hypothesis,
m = 3
(x1, y1) = (4, 6)
Plugging in the values in the formula and on further simplification, we get
y = 3x - 6
Hence, option (b) is your answer.